The intensity of light​ (in foot-candles) varies inversely as the square of​ x, the distance in feet from the light source. The intensity of light 2 feet from the source is 80 ​foot-candles. How far away is the source if the intensity of light is 5 ​foot-candles?


1
Expert's answer
2022-04-12T14:22:58-0400

Let "A" be the numerator of our formula, and "I" - intensity of light, then: "I=\frac{A}{X^2}"

If "X=2:80=\frac{A}4\implies A=320;"

If "I=5:5=\frac{320}{X^2}\implies X^2=\frac{320}5=64\implies X=8" ​foot-candles.